evaluate without using table 3 sin^2 45 + 2 cos^2 60

Grade:11

1 Answers

Pranay
6 years ago
3sin^2 45 + 2cos^2 60,As sin45 =1/√2, sin45^2=1/2,Similarly cos^260=1/4Si,3/2+ 1/2= 2Hope this would help